OpenStudy (anonymous):

Change the Cartesian double integral ∫1 −1∫−√1−x2 √1−x2 2 (1+x2+y2)2 dydx into an equivalent polar integral and evaluate. Change the Cartesian double integral ∫1 −1∫−√1−x2 √1−x2 2 (1+x2+y2)2 dydx into an equivalent polar integral and evaluate. @Mathematics

OpenStudy (unklerhaukus):

is this your integral? \[\int\limits_{1}^{-1}\int\limits_{-\sqrt{(1-x^2)}}^{2} (1+x^2+y^2)^2dxdy\]

OpenStudy (anonymous):

first integral is right the second should be \[\int\limits_{-\sqrt{1-x^2)}}^{\sqrt{1-x^2}}\] then the 2 is over the (1+x^2+y^2)^2

OpenStudy (unklerhaukus):

\[\int\limits_{-1}^{1}\int\limits_{-\sqrt{1-x^2}}^{\sqrt{1-x^2}} {2 \over (1+x^2+y^2)^2} dxdy \]